OpenStudy (anonymous):

Find derivative, quotient rule: (t+sec(t))/t^3

OpenStudy (shamil98):

You could also just use the product rule . (t+sec(t)) t^-3

OpenStudy (anonymous):

\[g(t)=\frac{ t +sect }{ t^3 }\]

OpenStudy (luigi0210):

Do you know the quotient rule?

OpenStudy (anonymous):

\[(\frac{ f }{ g })'(x)=(\frac{ f'(x)g(x)-f(x)g'(x) }{ g^2 (x) })\]

OpenStudy (luigi0210):

So what do you need help with exactly?

OpenStudy (luigi0210):

Setting it up, finding the derivative of a certain function?

OpenStudy (anonymous):

Setting it up part I don't understand.

OpenStudy (luigi0210):

It's pretty much this: \[\LARGE \frac{(t+sect)'(t^2)-(t+sect)(t^2)'}{(t^2)^2}\]

OpenStudy (anonymous):

@Luigi0210 is that the first step?

OpenStudy (luigi0210):

Yup, now just differentiate and simply :)